The ratio of oranges to grapefruit is 3:2. If there ae 60 fruits, the number of oranges is

Answer by jim_thompson5910(35256)

Let x = # of oranges and y = # of grapefruit

x+y = 60 ---> y = 60-x


3/2 = x/y


3/2 = x/(60-x)


3(60-x) = 2x


180-3x = 2x


180 = 2x+3x


180 = 5x


180/5 = x


36 = x


x = 36


y = 60-x


y = 60-36


y = 24


So there are 36 oranges and 24 grapefruit

If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim